The mean cost of domestic airfares in the United States rose to an all-time high of $385 per ticket. Airfares were based on the total ticket value, which consisted of the price charged by the airlines plus any additional taxes and fees. Assume domestic airfares are normally distributed with a standard deviation of $120. Use Table 1 in Appendix B.

a. What is the probability that a domestic airfare is $530 or more (to 4 decimals)?

b. What is the probability that a domestic airfare is $265 or less (to 4 decimals)?

c. What if the probability that a domestic airfare is between $320 and $480 (to 4 decimals)?

d. What is the cost for the 5% highest domestic airfares? (rounded to nearest dollar)
$  or - Select your answer -morelessItem 5
Answer #1

Solution :- Given: - Mean W = $ 385 std devi (0) = $ 120. a) P(x> 530) = 1- PC X 530) T-px-< 530-385) 1- P12 121) = 1-0.8869d) The a distribution of the highest s% is, P(Z > 2) = 5 % P( 272) = 0.05. 1 - P(Z <2) = 0.05. P(2<z) = 1-0.05. P(Z <1.645) =